# A Heterogeneous Dynamic Convolutional Neural Network for Image Super-resolution

### v2

@@ Abstract
Convolutional neural networks can automatically learn features via deep network architectures and given input samples. However, robustness of obtained models may have challenges in varying scenes. Bigger differences of a network architecture are beneficial to extract more complementary structural information to enhance robustness of an obtained super-resolution model. In this paper, we present a heterogeneous dynamic convolutional network in image super-resolution (HDSRNet). To capture more information, HDSRNet is implemented by a heterogeneous parallel network. The upper network can facilitate more contexture information via stacked heterogeneous blocks to improve effects of image super-resolution. Each heterogeneous block is composed of a combination of a dilated, dynamic, common convolutional layers, ReLU and residual learning operation. It can not only adaptively adjust parameters, according to different inputs, but also prevent long-term dependency problem. The lower network utilizes a symmetric architecture to enhance relations of different layers to mine more structural information, which is complementary with a upper network for image super-resolution. The relevant experimental results show that the proposed HDSRNet is effective to deal with image resolving. The code of HDSRNet can be obtained at https://github.com/hellloxiaotian/HDSRNet.
@@ I Introduction
Single image super-resolution (SISR) techniques can obtain high-quality images from given low-resolution images (LR), according to solution of ill-posed inverse problem [1, 2]. In recent years, machine learning techniques have obtained huge achievements in many applications, i.e., disease diagnosis [3], classification [4], object recognition [5], multimodal data fusion [6, 7], Metaverse [8, 9] and SISR [10]. Specifically, deep learning techniques with end-to-end architectures have obtained higher performance in image super-resolution. Deep convolutional neural networks (CNNs) use end-to-end architectures rather than manual setting parameters to obtain strong learning abilities to improve visual effects of image super-resolution (SR) [11]. For instance, Dong et al. designed 3-layer CNN in pixel mapping manner to convert a low-resolution (LR) image to a high-resolution (HR) image [12]. Although it can improve resolution of predicted images, it cannot make a tradeoff between network depth and performance in image super-resolution. To address this issue, scholars try to enlarge network depth to pursue improved performance in image super-resolution [13]. Stacking some small filters is used to achieve a very deep SR network [13]. Also, a residual learning technique is used on a deep layer of a deep network to make a balance between SR performance and computational costs. A deeply-recursive convolutional network (DRCN) [14] and deep recursive residual network (DRRN) [15] can exploit recursive networks to decrease parameters of training a SR model. To reduce computational complexities, some up-sampling operations, i.e., a transposed convolution [16], deconvolution and a sub-pixel convolution [17] are set on a deep layer to amplify low-frequency features for constructing HR images. For instance, a fast SR convolutional neural network (FSRCNN) [18] directly inputted low-resolution images to a deep CNN to obtain low-frequency features and applied a sub-pixel convolutional layer to transform low-frequency features to high-frequency features and obtain HR images. To further improve SR performance, an enhanced deep SR network as well as EDSR used enhanced residual blocks to extract more structural information for image super-resolution [19]. Also, deleting many unnecessary batch normalization can improve effect of training efficiency a SR model [19]. Although these algorithms are effective for SISR, they still may suffer from challenges for complex scenes.
In this work, we propose a heterogeneous dynamic convolutional network for SISR as well as HDSRNet. HDSRNet uses a heterogeneous parallel network to capture more information to improve performance of image SR. The upper network uses stacked heterogeneous blocks to extract more contexture information in image SR. Each heterogeneous block includes a dilated, dynamic, common convolutional layer, ReLU and residual learning operation to adaptively adjust parameters, according to different inputs. Also, it can prevent long-term dependency problem. The lower network can use a symmetric architecture to enhance relationships of different layers to obtain more complementary structural information. Quantitative and qualitative analysis show that the proposed method is a good tool for image super-resolution.
@@
(1) A heterogeneous parallel network is used to facilitate complementary structural information to improve performance for image super-resolution in terms of contextual and hierarchical information.
(2) Dynamic convolutions are embedded into a convolutional neural network to enhance robustness of obtained super-resolution model for complex scenes.
@@
The remaining of this paper is organized as follows. Section II describes related work. Section III illustrates the proposed method. Section IV collects massive comparative experimental results. And Section V summarizes our work.
@@ II-A Deep CNNs for Image Super-resolution
Due to shooting distance, captured images by cameras are unclear. Also, traditional image super-resolution methods need manual setting parameters and complex optimization parameters. To overcome this issue, deep learning techniques are extended for SISR [20, 21]. For instance, a deep recursive residual network used local residual connection and unit to improve effects of image super-resolution [15]. Alternatively, Yang et al. used recursive and gate units to transfer information shallow layers to deep layers to address long-term dependency problem [22]. To improve efficiency of image SR, given LR images are directly as an input of a convolutional neural network and high-quality images are constructed via an up-sampling operation in deep layer [18]. For instance, Ahn et al. used group convolutions to implement a cascading residual CNN to extract more robustness low-frequency information and decrease the number of parameters without causing significantly performance loss in SISR [23]. Tian et al. exploited symmetric group convolutional block to enhance relations of different channels to mine more accurate low-frequency information for image super-resolution [24]. Zhang et al. extended the depth of network and fully used hierarchical information to facilitate more low-frequency information for image SR [25]. According to mentioned methods, we can see that deep CNNs are good tools to obtain clearer images. Motivated by that, we propose a deep CNN for image super-resolution.
@@ II-B Dynamic convolution
Existing CNNs can share parameters to extract useful features to better represent images. However, they may suffer from challenges from varying input images in complex scenes [26]. To address this problem, dynamic convolutions are presented [26]. That is, they can dynamically adjust parameters to adaptively learn models for image applications, according to different inputs. For instance, Wan et al. achieved a graph convolution by arbitrarily structuring non Euclidean data for hyperspectral image classification [27]. Alternatively, Ding et al. can atomically find valuable receptive field of each target node to adaptively capture neighbor information in hyperspectral image classification [28]. To deal with native effects from incurring artifacts of textureless and edge regions, Duan et al. can dynamically generate kernels by region context from input images to better achieve image fusion [29]. To adaptively adjust brightness of enhanced images, Dai et al. fused Taylor expansion and dynamic convolution into a Retinex to intelligently improve clarity of low-light images and degree of brightness flexibly [30]. To extract more context information, Hou et al. proposed dynamic hybrid gradient convolution and coordinate sensitive attention to enhance boundary information extraction for remote sensing image segmentation [31]. To find more high-frequency information, Shen et al. used a spatially enhanced kernel generation to dynamic learn high-frequency and multi-scale features to better achieve denoising effect [32]. Additionally, Tian et al. combined dynamic convolution and wavelet transform to train an adaptive denoiser, according to different given noisy images [33]. According to illustrations, we find that dynamic convolutions are effective for image applications. Motivated by that, we use dynamic convolutions in a CNN for image super-resolution in this work.
@@ III-A Network Architecture
The proposed 18-layer HDSRNet contains two 16-layer parallel heterogeneous networks and a 2-layer construction block. The 16-layer parallel heterogeneous is composed of a 16-layer heterogeneous upper network and symmetrical lower network. The heterogeneous upper network is composed of a Conv+ReLU and five stacked heterogeneous blocks, which can extract more contexture information for image super-resolution. A Conv+ReLU is a composite of a convolutional layer and a ReLU operation, which is used to extract non-linear information from given low-resolution images. Also, its input and output channel numbers are 3 and 64, respectively. Its kernel size is $3\times 3$.These stacked heterogeneous blocks utilize different convolutional layers (i.e., dynamic and common convolutional layers) and ReLU to dynamically adjust parameters to obtain robust low-frequency information, according to low-resolution images of different inputs. To obtain complementary low-frequency information, a 16-layer symmetrical lower network is designed. It depends on a symmetrical architecture to enhance inter hierarchical connections to obtain more complementary structural information. Also, two sub-networks can interact information via a multiplication operation. A 2-layer construction block is used to transform low-frequency information to high-frequency information and construct predicted high-quality images. The process can be illustrated via the following formulates.
$\begin{array}[]{l}{I_{S}}=HDSRNet({I_{L}})\\ {\rm{}\quad}=CB{\rm{(}}HUNet({I_{L}})\times SLNet({I_{L}}))\\ {\rm{}\quad}=CB{\rm{((}}5HB{\rm{(}}CR({I_{L}})){\rm{)}}\times SLNet({I_{L}}))\\ {\rm{}\quad}=CB({O_{HUNet}}\times{O_{SLNet}})\end{array}$ (1)
where $HDSRNet$, $HUNet$ and $SLNet$ denote functions of HDSRNet, heterogeneous upper network and symmetrical lower network, respectively. $CR$ is a Conv+ReLU. $5HB$ is five stacked heterogeneous blocks. $CB$ denotes a construction block. $\times$ represents multiplication operation. Parameters of obtained HDSRNet can be obtained via a loss function in Section III. B. ${O_{HUNet}}$ and ${O_{SLNet}}$ are outputs of the HUNet and SLNet, respectively.
Figure 1: Network architecture of the proposed HDSRNet.
@@ III-B Loss Function
HDSRNet choose mean absolute error (MAE) [34] as the loss function to obtain parameters. Work process of MAE in the HDSRNet can be represented as below.
$l(p)={1\mathord{\left/{\vphantom{1T}}\right.\kern-1.2pt}T}\sum\limits_{j=1}^{T}{\left|{HDSRNet(I_{L}^{j})-I_{H}^{j}}\right|}$ (2)
where $I_{L}^{j}$ and $I_{H}^{j}$ denote the $jth$ low- and high-resolution images, respectively. $T$ represents the number of low-resolution images. $l$ is a loss function of HDSRNet. Also, $p$ stands for parameters of the trained HDSRNet. Also, it can be optimized by the Adam optimizer [35].
@@ III-C Heterogeneous Block
To train a robust denoiser, heterogeneous blocks are conducted to dynamically adjust parameters to obtain robust low-frequency information, according to different input low-resolution images. Each heterogeneous block is composed of a dilated Conv+ReLU, dynamic Conv+ReLU and Conv+ReLU, where dilated Conv+ReLU represents a combination of a dilated convolution [36] and a ReLU [37]. That is used to capture more context information. A dynamic Conv+ReLU is a combination of a dynamic convolution [26] and ReLU, where can adaptively learn parameters, according to different input information. To prevent long-term dependency, a residual operation is acted between an input of a dilated Conv+ReLU and output of Conv+ReLU. All the convolutional kernels are $3\times 3$. Input, output channel numbers, i.e., dilated, dynamic and common convolutional layers are 64. Also, dilated factor is 2 in the dilated convolutional layers. The mentioned process can be conducted the following equation.
$HB({O_{t}})=CR(DyCR(DiCR({O_{t}})))+{O_{t}}$ (3)
where ${O_{t}}$ is used to define an input of a heterogeneous block. $DiCR$ stands for a dilated Conv+ReLU. $DyCR$ represents a dynamic Conv+ReLU. $+$ is used to express a residual learning operation as well as $\oplus$ in Fig. 1.
@@ III-D Symmetrical Lower Sub-network
To obtain complementary low-frequency information, a 16-layer symmetrical lower network is conducted. Each layer contains a Conv+ReLU, where its input and output channel number are 64 besides the first layer, its kernel is $3\times 3$. Input and output channel number of the first layer are 3 and 64, respectively. To enhance relations of different layers, residual learning operations are used to act between the 1st and 16th, the 2nd and 15th, 3nd and 14th, 4th and 13th, 5th and 12th, 6th and 11th, 7th and 10th, 8th and 9th layers to transfer obtained information of shallow layers to deep layers to prevent long-term dependency and obtain robust information for image super-resolution. The procedure can be summarized as Eq. (4).
$\begin{array}[]{c}{O_{SLNet}}=CR(CR(...(CR(CR({O_{8}}+CR({O_{8}}))\\ \qquad+{O_{7}})+{O_{6}})+....)+{O_{2}})+{O_{1}}\end{array}$ (4)
where ${O_{i}}$ denotes an output of the $ith$ layer and $i=1,2,3...,8$. Also, ${O_{i}}=iCR({I_{L}})$, where $iCR$ stands for $i$ stacked Conv+ReLU.
@@ III-E Construction Block
A 2-layer construction block is used to construct predicted HR images. It contains two phases. The first phase uses a sub-pixel convolutional layer to convert low-frequency information to high-frequency information, its input and output channel number are 128 and 64, respectively. The second phase only utilizes a convolutional layer (Conv) to construct predicted resolution images, where its input and output channel number are 64 and 3, respectively. Their kernel sizes are $3\times 3$. These illustrations can be symbolled as Eq. (5).
$\begin{array}[]{l}{O_{S}}=CB({O_{HUNet}}\times{O_{SLNet}})\\ \quad\quad=C(Sub({O_{HUNet}}\times{O_{SLNet}}))\end{array}$ (5)
where $Sub$ and $C$ are functions of a sub-pixel convolutional layer and a convolutional layer, respectively.

### v3

# Adaptive Convolutional Neural Network for Image Super-resolution

@@ Abstract
Convolutional neural networks can automatically learn features via deep network architectures and given input samples. However, the robustness of obtained models may face challenges in varying scenes. Bigger differences in network architecture are beneficial to extract more diversified structural information to strengthen the robustness of an obtained super-resolution model. In this paper, we proposed a adaptive convolutional neural network for image super-resolution (ADSRNet). To capture more information, ADSRNet is implemented by a heterogeneous parallel network. The upper network can enhance relation of context information, salient information relation of a kernel mapping and relations of shallow and deep layers to improve performance of image super-resolution. That can strengthen adaptability of an obtained super-resolution model for different scenes. The lower network utilizes a symmetric architecture to enhance relations of different layers to mine more structural information, which is complementary with a upper network for image super-resolution. The relevant experimental results show that the proposed ADSRNet is effective to deal with image resolving. Codes are obtained at https://github.com/hellloxiaotian/ADSRNet.
@@ I Introduction
Single image super-resolution (SISR) techniques can obtain high-quality images from given low-resolution images (LR), according to solution of ill-posed inverse problems [1, 2]. In recent years, machine learning techniques have obtained huge achievements in various applications, i.e., disease diagnosis[3], classification[4, 5], object recognition[6], multimodal data fusion[7, 8], Metaverse[9, 10] and SISR [11, 12]. Specifically, deep learning techniques with end-to-end architectures have obtained better visual effect in SISR [13]. Deep convolutional neural networks (DCNNs) use end-to-end architectures rather than manual setting parameters to obtain strong learning abilities to improve visual effects of SISR[14]. Dong et al. designed 3-layer network through pixel transform to convert a low-resolution (LR) image to a high-resolution (HR) image [15]. Although this method can improve resolution of predicted images, it is limited between network depth and performance in SISR. To address this issue, scholars try to enlarge network depth to pursue improved performance in image super-resolution [16]. Stacking small filters is used to achieve a very deep SR network [16]. Additionally, a residual learning technique (RL) is employed on a deep network layer in a network to make a balance between SR performance and computational costs. A deeply-recursive convolutional network (DRCN) [17] and deep recursive residual network (DRRN) [18] can exploit recursive networks to decrease parameters of training a SR model. To reduce computational complexities, some up-sampling operations, such as transposed convolution [19], deconvolution and sub-pixel convolution [20] are set on a deep network layer to amplify low-frequency information for constructing HR images. For instance, a fast SR convolutional neural network (FSRCNN) [21] directly inputted low-resolution images to a deep CNN to obtain low-frequency features and applies a upsampling operation to transform low-frequency features to high-frequency features and obtain HR images. To further improve SR performance, an enhanced deep SR network as well as EDSR used enhanced residual blocks to extract more structural information for image super-resolution [22]. Also, deleting many unnecessary batch normalization can improve effect of training efficiency a SR model [22]. Although these algorithms are effective for SISR, they did not automatically update parameters, according to different scenes. That may cause poor robustness of obtained SR models for complex scenes. we propose a heterogeneous dynamic network for SISR as well as ADSRNet. ADSRNet uses a heterogeneous parallel network to capture more information to improve performance of image SR in this paper. The upper network uses stacked heterogeneous blocks to extract more contexture information in image SR. Also, each heterogeneous block uses different components to achieve dynamically learning parameters to strengthen robustness of our ADSRNet for SISR, according to different scenes. Also, it can prevent long-term dependency problem. The lower network can use a symmetric architecture to enhance relationships of between different layers to obtain more complementary structural information. Quantitative and qualitative analysis show that the proposed method is a good tool for image super-resolution.
@@
(1) A heterogeneous parallel network is used to facilitate complementary structural information to improve performance for image super-resolution in terms of contextual, hierarchical and salient information.
(2) Dynamic convolutions are embedded into a convolutional neural network to strengthen robustness of obtained SISR model for complex scenes.
@@
The remaining of this paper is listed. Section II describes related work. Section III illustrates the proposed method. Section IV collects massive comparative experimental results. And Section V summarizes our work.
@@ II-A Deep CNNs for Image Super-resolution
Due to the shooting distance, captured images by cameras are unclear. Also, traditional image super-resolution methods need manual setting parameters and complex optimization parameters. To address this issue, deep learning techniques are extended for SISR [23, 24]. For instance, a deep recursive residual network used local residual connection and unit to improve effects of image super-resolution [18]. Alternatively, Yang et al. used recursive and gate units to transfer information shallow layers to deep layers to address long-term dependency problem [25]. To improve efficiency of image SR, given LR images are directly as an input of a convolutional neural network and high-quality images are constructed via an up-sampling operation in deep layer [21]. For instance, Ahn et al. used group convolutions to implement a cascading residual CNN to extract more robustness low-frequency information and decrease the number of parameters without causing significantly performance loss in SISR [26]. Tian et al. exploited symmetric network architecture to mine more accurate low-frequency information for image super-resolution [27]. Zhang et al. extended the depth of network and fully used hierarchical information to afford more low-frequency information for image SR [28]. According to mentioned methods, we can see that deep CNNs are good tools to obtain clearer images and enhanced CNNs are effective for SISR. Thus, we enhanced a CNN to pursue better visual effects in SISR. Differing from previous research, we choose parallel and serial ways to facilitate heterogeneous information to better represent low-frequency structural information to improve visual effects of SISR.
@@ II-B Dynamic convolution
Existing CNNs can share parameters to extract useful features for better represent images. However, they may face challenges from varying input images in complex scenes [29]. To address this issue, dynamic convolutions have been introduced [29]. That is, they can dynamically adjust parameters to adaptively learn models for image applications, according to different inputs. For instance, Wan et al. achieved a graph convolution by arbitrarily structuring non Euclidean data for hyperspectral image classification [30]. Alternatively, Ding et al. can atomically find valuable receptive fields of each target node to adaptively capture neighbor information in hyperspectral image classification [31]. To deal with native effects from incurring artifacts of textureless and edge regions, Duan et al. can dynamically generate kernels by region context from input images to better achieve image fusion [32]. To adaptively adjust brightness of enhanced images, Dai et al. fused Taylor expansion and dynamic convolution into a Retinex to intelligently improve clarity of low-light images and degree of brightness flexibly [33]. To extract more context information, Hou et al. proposed dynamic hybrid gradient convolution and coordinate sensitive attention to enhance boundary information extraction for remote sensing image segmentation [34]. To find more high-frequency information, Shen et al. used a spatially enhanced kernel generation to dynamic learn high-frequency and multi-scale features to better achieve visual effect in image restoration [35]. Additionally, Tian et al. exploited relations of channels to extract richer low-frequency information for SISR [27]. Based on these examples, dynamic convolutions can dynamically learn parameters to improve robustness of an obtained CNN for SISR, according to different inputs. Thus, we use dynamic convolutions to guide a CNN to update parameters to strengthen adaptability in terms of recovering high-quality images for different scenes. Differing form previous research, a heterogeneous network architecture can facilitate more robustness features to generate high-quality images.
@@ III-A Network Architecture
The proposed 18-layer ADSRNet contains two 16-layer parallel heterogeneous networks and a 2-layer construction block. The 16-layer parallel heterogeneous is composed of a 16-layer heterogeneous upper network and a symmetrical lower network. The heterogeneous upper network is composed of a Conv+ReLU (CRU) and five stacked heterogeneous blocks, which can extract more contexture information for image super-resolution. A CRU is a composite of a convolutional layer and a ReLU operation, which is used to extract non-linear information from given low-resolution images. Also, its input and output channel numbers are 3 and 64, respectively. Its kernel size is $3\times 3$.These stacked heterogeneous blocks utilize different convolutional layers (i.e., dynamic and common convolutional layers) and ReLU to dynamically adjust parameters to obtain robust low-frequency features, according to low-resolution images of different inputs. To obtain complementary low-frequency features, a 16-layer symmetrical lower network is designed. It depends on a symmetrical architecture to enhance inter hierarchical connections to obtain more complementary structural information. Also, two sub-networks can interact information via a multiplication operation. A 2-layer construction block is used to transform low-frequency features to high-frequency features and construct predicted high-quality images. The process can be illustrated via the following formulates.
$\begin{array}[]{l}{I_{S}}=ADSRNet({I_{L}})\\ {\rm{}\quad}=CB{\rm{(}}HUNet({I_{L}})\times SLNet({I_{L}}))\\ {\rm{}\quad}=CB{\rm{((}}5HB{\rm{(}}CR({I_{L}})){\rm{)}}\times SLNet({I_{L}}))\\ {\rm{}\quad}=CB({O_{HUNet}}\times{O_{SLNet}}),\end{array}$ (1)
where $ADSRNet$, $HUNet$ and $SLNet$ are used as ADSRNet, heterogeneous upper network and symmetrical lower network, respectively. $CR$ is a CRU. $5HB$ is five stacked heterogeneous blocks. $CB$ denotes a construction block. $\times$ represents multiplication operation. Parameters of obtained ADSRNet can be obtained via a loss function in Section III. B. ${O_{HUNet}}$ and ${O_{SLNet}}$ are outputs of the HUNet and SLNet, respectively. $I_{L}$ and $I_{S}$ represent the input low-resolution image and the output high-resolution image, respectively.
Figure 1: Network architecture of the proposed ADSRNet.
@@ III-B Loss Function
ADSRNet choose mean absolute error (MAE) [36] as the loss function to obtain parameters. Work process of MAE in the ADSRNet can be represented as below.
$l(p)={1\mathord{\left/{\vphantom{1T}}\right.\kern-1.2pt}T}\sum\limits_{j=1}^{T}{\left|{ADSRNet(I_{L}^{j})-I_{H}^{j}}\right|},$ (2)
where $I_{L}^{j}$ and $I_{H}^{j}$ denote the $jth$ low- and high-resolution images, respectively. $ADSRNet$ demote function of ADSRNet, respectively. $T$ represents the number of low-resolution images. $l$ is a loss function of ADSRNet. Also, $p$ stands for parameters of the trained ADSRNet. Also, it can be optimized by the Adam optimizer [37].
@@ III-C Heterogeneous Block
To improve the SISR effect and robustness of the network, heterogeneous blocks are conducted to dynamically adjust parameters to obtain robust low-frequency information, according to different input low-resolution images. Each heterogeneous block is composed of a dilated CRU, dynamic CRU and CRU, where dilated CRU represents a combination of a dilated convolution [38] and a ReLU [39]. That is used to capture more context information. A dynamic CRU is a combination of a dynamic convolution [29] and ReLU, where can adaptively learn parameters, according to different input information. To resolve long-term dependency, a RL operation is employed between an input of a dilated CRU and output of CRU. Mentioned convolutional kernels are $3\times 3$. Input channel number and output channel number, i.e., dilated, dynamic and common convolutional layers are 64. Also, dilated factor is 2 in the dilated convolutional layers. The mentioned process can be conducted the following equation.
$HB({O_{t}})=CR(DyCR(DiCR({O_{t}})))+{O_{t}},$ (3)
where ${O_{t}}$ is used to define an input of a heterogeneous block. $CR$ is a CRU. $DiCR$ stands for a dilated CRU. $DyCR$ represents a dynamic CRU. $+$ is used to express a RL operation also regarded to $\oplus$ in Fig. 1.
@@ III-D Symmetrical Lower Sub-network
To obtain complementary low-frequency information, a 16-layer symmetrical lower network is conducted. Each layer contains a CRU, where its channel number of input and output are 64 besides the first layer, its kernel is $3\times 3$. Channel number of input and output from the first layer are set to 3 and 64, respectively. To enhance relations of different layers, residual learning operations are used to act between the 1st and 16th, the 2nd and 15th, 3nd and 14th, 4th and 13th, 5th and 12th, 6th and 11th, 7th and 10th, 8th and 9th layers to transfer obtained information of shallow layers to deep layers to prevent long-term dependency and obtain robust information for SISR. The procedure can be summarized as Eq. (4).
$\begin{array}[]{c}{O_{SLNet}}=CR(CR(...(CR(CR({O_{8}}+CR({O_{8}}))\\ \qquad+{O_{7}})+{O_{6}})+....)+{O_{2}})+{O_{1}},\end{array}$ (4)
where ${O_{i}}$ denotes an output of the $ith$ layer and $i=1,2,3...,8$. Also, ${O_{i}}=iCR({I_{L}})$, where $iCR$ stands for $i$ stacked CRU. $O_{SLNet}$ represent the output of SLNet, respectively.
@@ III-E Construction Block
A 2-layer construction block is used to construct predicted HR images. It includes two steps. The first step is acted to a sub-pixel convolutional layer, which can use low-frequency features to obtain high-frequency features. Its input and output channel number are 128 and 64, respectively. The second phase only utilizes a convolutional layer (Conv) to construct predicted resolution images, where its input and output channel number are 64 and 3, respectively. Their kernel sizes are $3\times 3$. These illustrations can be symbolled as Eq. (5).
$\begin{array}[]{l}{O_{S}}=CB({O_{HUNet}}\times{O_{SLNet}})\\ \quad\quad=C(Sub({O_{HUNet}}\times{O_{SLNet}})),\end{array}$ (5)
where $Sub$ and $C$ are functions of a sub-pixel convolutional layer and a convolutional layer, respectively. $CB$ denotes a construction block. $O_{HUNet}$ and $O_{SLNet}$ express the output of HUNet and SLNet.